Imagine that you made a purchase that cost $1.85. You paid with a $10.00 bill, and you received $8.15 in change. Notice that the digits in your change are the same as the digits in the amount you spent. Using a $10.00 bill as payment, there are other amounts you cold spend that would result in change that has the same digits as the amount you spent. What are these combinations? How many can you find? How did you find them?
Answer by checkley71(8403)

OTHER COMBINATIONS ARE:
0.95+9.05=10.00
2.75+7.25=10.00
3.65+6.35=10.00
4.55+5.45=10.00
5.45+4.55=10.00
6.35+3.65=10.00
7.25+2.75=10.00
8.15+1.85=10.00
9.05+0.95=10.00
9.05+0.95=10.00